OpenStudy (anonymous):

3a-2(b+2) when a = 5 and b= -4

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

You are missing an important word in your question. It should read Evaluate 3a-2(b+2) when a = 5 and b= -4

OpenStudy (skullpatrol):

Evaluating a variable expression like 3a-2(b+2) when a = 5 and b= -4 means replacing each variable, a and b, in the expression: 3a-2(b+2) by the given value: a = 5 and b= -4 and simplifying the result.

OpenStudy (skullpatrol):



OpenStudy (anonymous):

Iv got the answer as 19 but im not sure if that is right?

OpenStudy (skullpatrol):

19 is correct :)